# tSAPPHIRES collaboration

### v1

@@ VI Conclusion
We have presented the results of a single-point ALP search using a variable-angle three-beam stimulated resonant photon collider (tSRPC). The energies of the creation lasers $c_{1}$ and $c_{2}$ ($1.4~\mathrm{mJ}/34~\mathrm{fs}$, Ti:Sapphire) and the inducing laser i ($60.6~\mathrm{mJ}/7.1~\mathrm{ns}$, Nd:YAG) were each increased by more than three orders of magnitude compared to the previous pilot search [25]. The experiment was performed in vacuum to suppress background photons. No statistically significant signal was observed. We therefore set exclusion limits, reaching a minimum coupling of $g/M=4.0\times 10^{-10}~\mathrm{GeV}^{-1}$ at $m_{a}=2.27~\mathrm{eV}$, based on the formulation dedicated to tSRPC [24] and the experimentally determined overlap factor presented in this work. This is the first model-independent upper limit that reaches the KSVZ benchmark in the eV range. This single-point search demonstrates the feasibility of a smooth mass scan by varying the collision angle in near future searches.
@@ Appendix A PMT calibration
In this measurement, we quantitatively evaluate single-photon signals with a photomultiplier tube (PMT). The photocathode quantum efficiency $\epsilon_{QE}$ is wavelength dependent, and the electron multiplication in the dynode chain is well described by a gamma distribution. We thus calibrated the one-photon-equivalent charge (1 PEQ) near the signal wavelength, using a pulsed white-light source (SuperK Compact; pulse width $t<1~\mathrm{ns}$) and a band-pass filter ($660\pm 5~\mathrm{nm}$).
@@
$G_{0}\!\left(q\,\big|\,m_{0},s_{0}\right)$ $=\frac{1}{\sqrt{2\pi}\,s_{0}}\exp\!\left[-\frac{(q-m_{0})^{2}}{2s_{0}^{2}}\right],$
$g_{n}\!\left(q\,\big|\,n\kappa,\theta\right)$ $=\frac{1}{\Gamma(n\kappa)\,\theta^{\,n\kappa}}\,(-q)^{\,n\kappa-1}\exp\!\left(\frac{q}{\theta}\right)\qquad(q\leq 0),$ (A.1)
@@
$\kappa=\left(\frac{m_{1}}{s_{1}}\right)^{\!2},\qquad\theta=-\frac{s_{1}^{2}}{m_{1}}.$
@@
$F(q;\lambda)=A\Bigg[e^{-\lambda}G_{0}\!\left(q\,\big|\,m_{0},s_{0}\right)+\sum_{n=1}^{\infty}\frac{e^{-\lambda}\lambda^{n}}{n!}\;g_{n}\!\left(q\,\big|\,n\kappa,\theta\right)\Bigg].$
@@
$Q(\lambda)=m_{1}\,\lambda+b.$
Figure 10: Linear relation between the fit-averaged number of photoelectrons and the mean charges. The linear fit is performed for data with $\lambda<10$, and the slope corresponds to the one-photon equivalent charge.
Figure 10 shows the mean value of charge distributions $Q$ versus photon expectation value $\lambda$ ($\lambda<10$) together with the best-fit line. The slope equals the one-photon-equivalent charge,
$1\,\mathrm{PEQ}\;\equiv\;m_{1}.$
Appendix B 3-beam overlap factor $\overline{\mathcal{D}}_{\mathrm{three}}$ including offset
@@
$\quad\times\exp\!\biggl[-2\biggl\{\sum_{j}\!\bigl(\xi_{j}^{2}+\eta_{j}^{2}+\zeta_{j}^{2}\bigr)-\frac{\Bigl(\sum_{j}(\mu_{j}\xi_{j}+\nu_{j}\zeta_{j})\Bigr)^{2}}{\sum_{j}(\mu_{j}^{2}+\nu_{j}^{2})}\biggr\}\biggr],$ (B.1)
@@
$\overline{\mathcal{D}}^{(r)}\;=\;\tfrac{1}{2}\,\bigl(\mathcal{D}_{0}+\mathcal{D}_{r}\bigr).$
@@
$\overline{\mathcal{D}}_{\mathrm{three}}\;=\;\frac{1}{N_{\mathrm{run}}}\sum_{r=1}^{N_{\mathrm{run}}}\overline{\mathcal{D}}^{(r)}.$

### v2

@@ VI Conclusion
We have presented the results of a single-point ALP search using a variable-angle three-beam stimulated resonant photon collider (tSRPC). The energies of the creation lasers $c_{1}$ and $c_{2}$ ($1.4~\mathrm{mJ}/34~\mathrm{fs}$, Ti:Sapphire) and the inducing laser i ($60.6~\mathrm{mJ}/7.1~\mathrm{ns}$, Nd:YAG) were each increased by more than three orders of magnitude compared to the previous pilot search [25]. The experiment was performed in vacuum to suppress background photons. No statistically significant signal was observed. We therefore set exclusion limits, reaching a minimum coupling of $g/M=4.2\times 10^{-10}~\mathrm{GeV}^{-1}$ at $m_{a}=2.27~\mathrm{eV}$, based on the formulation dedicated to tSRPC [24] and the experimentally determined overlap factor presented in this work. This is the first model-independent upper limit that reaches the KSVZ benchmark in the eV range. This single-point search demonstrates the feasibility of a smooth mass scan by varying the collision angle in near future searches.
@@ Appendix A PMT calibration
In this measurement, we quantitatively evaluate single-photon signals with a photomultiplier tube (PMT). The photocathode quantum efficiency $\epsilon_{QE}$ is wavelength dependent, and the electron multiplication in the dynode chain is well described by a gamma distribution. We thus calibrated the one-photon-equivalent charge near the signal wavelength, using a pulsed white-light source (SuperK Compact; pulse width $t<1~\mathrm{ns}$) and a band-pass filter ($660\pm 5~\mathrm{nm}$).
@@
$G_{0}\!\left(q\,\big|\,m_{0},s_{0}\right)$ $=\frac{1}{\sqrt{2\pi}\,s_{0}}\exp\!\left[-\frac{(q-m_{0})^{2}}{2s_{0}^{2}}\right],$ (23)
$g_{n}\!\left(q\,\big|\,n\kappa,\theta\right)$ $=\frac{1}{\Gamma(n\kappa)\,\theta^{\,n\kappa}}\,(-q)^{\,n\kappa-1}\exp\!\left(\frac{q}{\theta}\right)\qquad(q\leq 0),$ (24)
@@
$\kappa=\left(\frac{m_{1}}{s_{1}}\right)^{\!2},\qquad\theta=-\frac{s_{1}^{2}}{m_{1}}.$ (25)
@@
$F(q;\lambda)=A\Bigg[e^{-\lambda}G_{0}\!\left(q\,\big|\,m_{0},s_{0}\right)+\sum_{n=1}^{\infty}\frac{e^{-\lambda}\lambda^{n}}{n!}\;g_{n}\!\left(q\,\big|\,n\kappa,\theta\right)\Bigg].$ (26)
@@
$Q(\lambda)=Q_{1PE}\,\lambda+b.$ (27)
Figure 11: Linear relation between the fit-averaged number of photoelectrons and the mean charges. The blue dots show the measurements in Fig. 10 (b)-(h) plus additional datasets not shown in Fig. 10. The linear fit is performed for data with $\lambda<10$, and the slope corresponds to the one-photon equivalent charge.
Figure 11 shows the mean value of charge distributions $Q$ versus photon expectation value $\lambda$ ($\lambda<10$) together with the best-fit line. The slope parameter $Q_{1PE}$ equals the one-photon-equivalent charge.
Appendix B Statistical uncertainty from Gamma-distributed single-photon charge fluctuations
In this appendix we derive the statistical uncertainty of the photon-number estimator used in the analysis, starting from the fact that the photomultiplier tube (PMT) records analog integrated charges rather than integer photon counts. The estimator is constructed from the measured charge using the calibrated one-photon-equivalent charge, and its variance receives contributions both from Poisson photon statistics and from the finite width of the single-photon charge response.
For clarity, we first consider the idealized case in which the photon number is fixed to a particular integer value $n$ and does not fluctuate from trigger to trigger. In that case the total charge can be written as
$Q_{n}=\sum_{j=1}^{n}q_{j},$ (28)
where $q_{j}$ denotes the charge contributed by the $j$-th detected photon ($j=1,\dots,n$). Each $q_{j}$ is an independent and identically distributed random variable corresponding to the single-photon Gamma response (the $n=1$ component of Eq. (24)), with mean $m_{1}$ and variance $s_{1}^{2}$. We use $m_{1}$ as the one-photon-equivalent charge. In this fixed-$n$ ensemble one has
$\langle Q_{n}\rangle=n\,m_{1},\qquad\mathrm{Var}(Q_{n})=n\,s_{1}^{2}.$ (29)
In the actual experiment, however, the photon number $n$ fluctuates from trigger to trigger according to the Poisson law with mean $\lambda$, so that
$\langle n\rangle=\lambda,\qquad\mathrm{Var}(n)=\lambda.$ (30)
The total variance of the measured charge $Q$ can then be decomposed, according to the law of total variance, as
$\mathrm{Var}(Q)$ $=\big\langle\mathrm{Var}(Q_{n})\big\rangle+\mathrm{Var}\big(\langle Q_{n}\rangle\big)$ (31)
$=\big\langle ns_{1}^{2}\big\rangle+\mathrm{Var}(nm_{1})$
$=s_{1}^{2}\langle n\rangle+m_{1}^{2}\mathrm{Var}(n)$
$=\lambda s_{1}^{2}+\lambda m_{1}^{2}$
$=\lambda m_{1}^{2}\left(1+\frac{s_{1}^{2}}{m_{1}^{2}}\right).$
It is convenient to define an effective Fano factor directly for the charge variable $Q$ as
$F\equiv\frac{\mathrm{Var}(Q)}{m_{1}\,\langle Q\rangle}=1+\frac{s_{1}^{2}}{m_{1}^{2}}=1+\frac{1}{\kappa},$ (32)
where we used $\kappa=(m_{1}/s_{1})^{2}$ from Eq. (25). For the measured single-photon response we obtain $\kappa\simeq 2.81$, corresponding to
$F\simeq 1.36.$ (33)
Thus the charge variance scales as
$\mathrm{Var}(Q)=F\,m_{1}^{2}\,\lambda.$ (34)
In practice, the Poisson mean $\lambda$ is not known a priori and is estimated from the data. Defining the photon-number estimator by
$N=\frac{Q}{m_{1}},$ (35)
we have $\langle N\rangle=\lambda$, so that $N$ is an unbiased estimator of the mean photon number. In the following we therefore replace $\lambda$ by $N$ and use
$\mathrm{Var}(Q)\simeq F\,m_{1}^{2}\,N$ (36)
in the error propagation.
The one-photon-equivalent charge $Q_{1PE}\equiv m_{1}$ is known with a finite calibration uncertainty $\delta Q_{1PE}$ from Eq. (10). Treating $Q$ and $Q_{1PE}$ as independent quantities, the standard error propagation for
$N=\frac{Q}{Q_{1PE}}$ (37)
gives
$(\delta N)^{2}$ $=$ $\left(\frac{\partial N}{\partial Q}\right)^{2}(\delta Q)^{2}+\left(\frac{\partial N}{\partial Q_{1PE}}\right)^{2}(\delta Q_{1PE})^{2}$ (38)
$=$ $\frac{\mathrm{Var}(Q)}{Q_{1PE}^{2}}+N^{2}\left(\frac{\delta Q_{1PE}}{Q_{1PE}}\right)^{2}.$
Using $\mathrm{Var}(Q)=FQ_{1PE}^{2}N$, we obtain
$(\delta N)^{2}=FN+N^{2}\left(\frac{\delta Q_{1PE}}{Q_{1PE}}\right)^{2}.$ (39)
The first term represents the genuine statistical uncertainty of the photon-number estimator, while the second term, induced by the finite precision of the calibration of $Q_{1PE}$, is treated as a global calibration systematic. For the present data set, this second term evaluates to $\sim 3\times 10^{-3}$, which is negligible compared to the leading statistical term and is therefore ignored in the following.
For each trigger pattern $X\in\{S,C,I,P\}$, the reconstructed photon number $N_{X}$ follows
$\mathrm{Var}(N_{X})\simeq F\,N_{X}.$ (40)
The signal photon yield per shot is obtained from the linear combination
$n_{\mathrm{sig}}=N_{S}-N_{C}-N_{I}+N_{P},$ (41)
so that, neglecting correlations between different patterns, the corresponding statistical variance is
$\delta n_{\mathrm{sig}}^{2}=F\left(N_{S}+N_{C}+N_{I}+N_{P}\right).$ (42)
This expression is used to evaluate the statistical error quoted in the main text.
Appendix C 3-beam overlap factor $\overline{\mathcal{D}}_{\mathrm{three}}$ including offset
@@ Appendix B Statistical uncertainty from Gamma-distributed single-photon charge fluctuations
$\quad\times\exp\!\biggl[-2\biggl\{\sum_{j}\!\bigl(\xi_{j}^{2}+\eta_{j}^{2}+\zeta_{j}^{2}\bigr)-\frac{\Bigl(\sum_{j}(\mu_{j}\xi_{j}+\nu_{j}\zeta_{j})\Bigr)^{2}}{\sum_{j}(\mu_{j}^{2}+\nu_{j}^{2})}\biggr\}\biggr],$ (43)
@@
$\overline{\mathcal{D}}^{(r)}\;=\;\tfrac{1}{2}\,\bigl(\mathcal{D}_{0}+\mathcal{D}_{r}\bigr).$ (44)
@@
$\overline{\mathcal{D}}_{\mathrm{three}}\;=\;\frac{1}{N_{\mathrm{run}}}\sum_{r=1}^{N_{\mathrm{run}}}\overline{\mathcal{D}}^{(r)}.$ (45)